Translate and find value . If the product of three numbers is 48 and two of the numnbers are 3 and 8 what is third number?
Answer by Susmitha(21)

let the three numbers be a, b and c.
given two numbers are 3 and 8.
let a=3 and b=8, c=?
a*b*c=48
3*8*c=48
24*c=48
c=48/24
c=2
therefore 2 is the required third number.
so you want me to solve for x in the equation 6x+4=40
6x=40-4
6x=36
x=36/6
x=6